# Can Unconfident LLM Annotations Be Used for Confident Conclusions?

### v1

@@ 1 Introduction
Our task is statistical estimation of a quantity of interest. We evaluate our approach on five estimation tasks in three CSS settings (politeness, stance, and media bias) in terms of confidence interval coverage and effective sample size, which measures the increase in accuracy due to augmenting human with LLM annotations (Sec. 3.4). We find that naively treating LLM annotations as human data can lead to highly inaccurate estimates and poor coverage. At the same time, our method maintains the target coverage, while outperforming the baselines (defined in Sec. 3.3) in terms of the effective sample size. The latter is enabled partially by the fact that in all tested settings the confidence scores are reflective of LLM accuracy.
@@ 2.1 LLMs for Data Annotation Tasks
LLMs have shown great potential in handling text-annotation tasks without prior task-specific training, sometimes even outperforming crowd workers Gilardi et al. (2023). NLP, LLMs offer transformative opportunities for any discipline that relies on text as data. Fields such as psychology, political science, sociology, communications, and economics recognize this emerging technology’s potential to enhance simulation-based research Bail (2024), and facilitate tasks such as text analysis, concept induction Lam et al. (2024), and topic modeling Pham et al. (2024).
@@ 3.2 Confidence-Driven Inference
$\hat{\theta}^{\mathrm{conf}}=\operatorname*{arg\,min}_{\theta}\frac{1}{n}\sum_{i=1}^{n}\left(\lambda\hat{\ell}_{\theta,i}+(\ell_{\theta,i}-\lambda\hat{\ell}_{\theta,i})\frac{\xi_{i}}{\pi_{i}}\right),$ (2)
@@ 3.3 Baselines
Finally, we consider the naive baseline which treats LLM annotations as human annotations, applying the standard estimator to those annotations and naively forming a confidence interval. This baseline does not suffer from a budget constraint, since LLM annotations are assumed to be cheap and available for all $n$ instances, but it may be biased if the LLM produces biased outputs.
@@ 4.1 Estimation tasks
Stance on global warming Gain in eff. sample size ( 33.77 $\pm$ 25.06)( 17.84 $\pm$ 25.67)%
Coverage 88% 94% 0%
@@
News headlines ($n=2,300$) are agreeing, neutral, or disagreeing with the stance that global warming is a serious concern Luo et al. (2020). Stance annotations facilitate the study of linguistic differences between media supporting or rejecting global warming, which have important implications for communication and policy Hmielowski et al. (2014). In this task, we estimate $\theta^{*}$ corresponding to $O_{\mathrm{agreement}}$, the odds ratio of agreement given the presence of affirming devices such as “expert,” “proven,” “renowned,” and so on. Formally, denoting by $X_{\mathrm{affirm}}\in\{0,1\}$ the presence of an affirming device and $H_{\mathrm{agree}}\in\{0,1\}$ the annotation of agreement, we have
@@ 4.2 Evaluation
Our main evaluation is based on LLM annotations collected with GPT-4o; analogous results with GPT-3.5 can be found in App. B.1. Table 2 in App. A.3 lists prompt texts and parameters. To test LLM performance out of the box, all annotations are collected using zero-shot prompting. Overall, the confidence scores are calibrated with accuracy, but the annotations are only in moderate agreement with human annotations in all three settings (see App. A.3). This is aligned with our lack of assumption that the LLM annotations are good: we want to produce a valid confidence interval no matter the quality of the LLM annotations.
We report the two key metrics (effective sample size and coverage), for the three selected settings (the study of politeness, stance, and bias), where the task is to estimate the five target quantities $\beta_{\mathrm{hedge}}$, $\beta_{\mathrm{1pp}}$, $O_{\mathrm{agreement}}$, $p_{\mathrm{left}}$, and $p_{\mathrm{right}}$. Both metrics are estimated over 100 trials for varying $n_{\mathrm{human}}$, the budget for human annotations.
Our main findings are reported in Figure 2 and summarized in Table 1. Figure 2 also depicts the computed confidence intervals in five randomly chosen trials across the five target quantities, for the lowest value of $n_{\mathrm{human}}$ in the considered range.
@@
Similarly, it is informative to consider the necessary budget of human annotations $n_{\mathrm{human}}$ given a desired effective sample size $n_{\mathrm{effective}}$. To achieve a desired fixed effective sample size, the confidence-driven approach reduces the needed number of human annotations across all five target quantities. For instance, to achieve $n_{\mathrm{effective}}=1000$, only between around 250 ($\beta_{\mathrm{1pp}}$) and 750 ($p_{\mathrm{left}}$) human annotations are needed, thus reducing the number of human annotations needed to achieve equally accurate estimates by at least 25% for all tasks.
@@
Table 1 summarizes the gain in effective sample size for $n_{\mathrm{human}}=500$. Across the five tasks, the confidence-driven approach achieves a substantial gain in the effective sample size, providing at minimum around +30% gain (when estimating $p_{\mathrm{left}}$), going even over +300% (when estimating $\beta_{\mathrm{1pp}}$). Again, the confidence-driven approach achieves a higher gain than the non-adaptive approach for each task, which can even be negative (when estimating $\beta_{\mathrm{hedge}}$ and $\beta_{\mathrm{1pp}}$).
@@ 5 Discussion
Finally, given the growing interest in harnessing the capabilities of LLMs across disciplines, the accessibility of a method is an important consideration. Researchers can simply prompt the LLM for its confidence level via API access and leverage Confidence-Driven Inference to combine LLM confidence with LLM and human annotations to produce a valid statistical estimate. This approach can be applied to a wide range of tasks, across fields.
6 Limitations
The external validity of our findings is contingent upon two key assumptions: that the text instances are i.i.d. from a relevant distribution, and that the researcher has full control of the annotation process. The first may be violated if the distribution of texts shifts over time, and the collected instances are no longer representative of the current quantity of interest. For example, it is possible that relationships between linguistic devices and perceived politeness evolve over time. The second assumption may be violated in situations where certain annotations are difficult to obtain (e.g., for low-resource languages). Our approach may lead to inaccurate or misleading conclusions under either violation. We thus caution against generalizing to settings where text instances exhibit time-varying shifts or the researcher is not in control over the data collection process.
If the adaptive sampling probabilities $\pi_{i}$ are poorly chosen—potentially due to inaccurate verbalized confidence scores—the resulting estimates could have a higher mean squared error (MSE) than if uniform, non-adaptive sampling were used. This could even result in an estimate with a larger MSE than the human-only baseline (for sensitivity to miscalibration, see App B.2). However, by using power tuning, as detailed in Section 3.2, we ensure that incorporating LLM annotation into the estimation process does not hurt the estimate (i.e., does not increase the MSE) regardless of the sampling method used for human annotations (whether uniform or adaptive). That said, if the confidence scores are not reliable (i.e., they are poorly calibrated), the estimate may become too noisy, leading to overly wide confidence intervals.
@@ 6 Limitations
Additionally, while we treat human annotations as the gold standard in our study, we acknowledge that human annotations are biased, and that reasonable annotators can disagree. Future work could explore ways to account for variability and bias in human annotations.
Human annotations are often obtained through crowdsourcing, which may itself be influenced by LLMs, as crowd workers might use LLMs to increase productivity Veselovsky et al. (2023). Although we use datasets collected before the widespread availability of LLMs, detecting AI-generated text remains a challenge Verma et al. (2024); Weber-Wulff et al. (2023).
@@ A.3 LLM and Human Annotation Details
Stage 1 GPT-4o annotations are in moderate agreement with human annotations in all three settings: $\kappa_{\mathrm{politeness}}=0.39$, $\kappa_{\mathrm{stance}}=0.57$, and $\kappa_{\mathrm{bias}}=0.43$.
@@ A.4 Computation of Evaluation Metrics
Appendix B Supplementary Results
B.1 An Alternative LLM
Figure 4 and Table 3 summarize the results using GPT-3.5. For $n_{\mathrm{human}}=500$, in each estimation task the confidence-driven approach again achieves a higher gain in effective sample size than the non-adaptive approach. Moreover, it always achieves a positive gain. In contrast, the non-adaptive approach achieves a negative gain in three out of the five estimation tasks (both politeness estimates and the stance estimate). The confidence-driven and non-adaptive approaches always achieve over 90% coverage. In contrast, LLM-only coverage is always poor using GPT-3.5 (while using GPT-4o it was poor on four out of the five estimation tasks).
B.2 Sensitivity to Miscalibrated LLM Confidence Scores
@@
Stance on global warming Gain in eff. sample size ( 24.66 $\pm$ 16.24)( -12.13 $\pm$ 16.95)%
Coverage 92% 94% 0%

### v2

@@ 1 Introduction
Our task is statistical estimation of a quantity of interest. We evaluate our approach on five estimation tasks in three CSS settings (politeness, stance, and media bias) in terms of confidence interval coverage and effective sample size, which measures the increase in accuracy due to augmenting human with LLM annotations (Sec. 3.4). We find that naively treating LLM annotations as human data can lead to highly inaccurate estimates and poor coverage. By contrast, our method maintains the target coverage, while outperforming the baselines (defined in Sec. 3.3) in terms of the effective sample size. The latter is enabled partially by the fact that, in all tested settings, the verbalized confidence scores reflect LLM accuracy. Higher confidence scores correspond to higher accuracy with respect to human annotations, allowing for a strategic selection of a smaller number of human annotations.
@@ 2.1 LLMs for Data Annotation Tasks
LLMs have shown great potential in handling text-annotation tasks without prior task-specific training, sometimes even outperforming crowd workers Gilardi et al. (2023); Zheng et al. (2023); Liu et al. (2023); Chiang and Lee (2023); Kim et al. (2023). NLP, LLMs offer transformative opportunities for any discipline that relies on text as data. Fields such as psychology, political science, sociology, communications, and economics recognize this emerging technology’s potential to enhance simulation-based research Bail (2024), and facilitate tasks such as text analysis, concept induction Lam et al. (2024), and topic modeling Pham et al. (2024).
@@ 3.2 Confidence-Driven Inference
[…]
@@ 3.3 Baselines
Finally, we consider the naive baseline which treats LLM annotations as human annotations, applying the standard estimator to those annotations and naively forming a confidence interval. This baseline does not suffer from a budget constraint, since LLM annotations are assumed to be cheap and available for all $n$ instances, but it may be biased.
@@ 4.1 Estimation tasks
Stance on global warming Gain in eff. sample size ( 102.51 $\pm$ 13.01)( 13.72 $\pm$ 22.79)%
Coverage 96% 96% 0%
@@
News headlines ($n=2,300$) are agreeing, neutral, or disagreeing with the stance that global warming is a serious concern Luo et al. (2020). Stance annotations facilitate the study of linguistic differences between media supporting or rejecting global warming, which have implications for communication and policy Hmielowski et al. (2014). In this task, we estimate $\theta^{*}$ corresponding to $O_{\mathrm{agreement}}$, the odds ratio of agreement given the presence of affirming devices such as “expert,” “proven,” “renowned,” and so on. Formally, denoting by $X_{\mathrm{affirm}}\in\{0,1\}$ the presence of an affirming device and $H_{\mathrm{agree}}\in\{0,1\}$ the annotation of agreement, we have
@@ 4.2 Evaluation
Our main evaluation is based on LLM annotations collected with GPT-4o; analogous results with GPT-3.5 can be found in App. B.1. Table 2 in App. A.3 lists prompt texts and parameters. To test LLM performance out of the box, all annotations are collected using zero-shot prompting. Our evaluation is designed to reflect a typical CSS use-case by using standard classification CSS tasks and testing popular API-based models, without any task-specific tuning or training. Analogous results with different formulations of the annotation task, prompting mechanisms, and models are outlined in App. B.
Overall, the confidence scores are calibrated with accuracy, but the annotations are only in moderate agreement with human annotations in all three settings (see App. A.3). This is aligned with our lack of assumption that the LLM annotations are good: we want to produce a valid confidence interval no matter the quality of the LLM annotations.
We report the two key metrics (effective sample size and coverage), for the three selected settings (the study of politeness, stance, and bias), where the task is to estimate the five target quantities $\beta_{\mathrm{hedge}}$, $\beta_{\mathrm{1pp}}$, $O_{\mathrm{agreement}}$, $p_{\mathrm{left}}$, and $p_{\mathrm{right}}$. Both metrics are estimated over 100 trials for varying $n_{\mathrm{human}}$, the budget for human annotations. Our main findings are reported in Figure 2 and summarized in Table 1.
@@
Similarly, it is informative to consider the necessary budget of human annotations $n_{\mathrm{human}}$ given a desired effective sample size $n_{\mathrm{effective}}$. instance, to achieve $n_{\mathrm{effective}}=1000$, only between around 250 ($\beta_{\mathrm{1pp}}$) and 750 ($p_{\mathrm{left}}$) human annotations are needed. We thus reduce the number of human annotations needed to achieve equally accurate estimates by at least 25% for all tasks.
@@
Table 1 summarizes the gain in effective sample size for $n_{\mathrm{human}}=500$. Across the five tasks, the confidence-driven approach achieves a substantial gain in the effective sample size, providing at minimum around +30% gain (when estimating $p_{\mathrm{left}}$), going even over +300% (when estimating $\beta_{\mathrm{1pp}}$). Again, the confidence-driven approach achieves a higher gain than the non-adaptive approach for each task, which can even be negative.
@@ 5 Discussion
Wer note that the external validity of our findings is contingent upon two key assumptions: that the text instances are i.i.d. from a relevant distribution, and that the researcher has full control of the annotation process. The first may be violated if the distribution of texts shifts over time, and the collected instances are no longer representative of the current quantity of interest. For example, it is possible that relationships between linguistic devices and perceived politeness evolve over time. The second assumption may be violated in situations where certain annotations are difficult to obtain (e.g., for low-resource languages). Our approach may lead to inaccurate or misleading conclusions under either violation. We thus caution against generalizing to settings where text instances exhibit time-varying shifts or the researcher is not in control over the data collection process.
If the adaptive sampling probabilities $\pi_{i}$ are poorly chosen—potentially due to inaccurate verbalized confidence scores—the resulting estimates could have a higher mean squared error (MSE) than if uniform, non-adaptive sampling were used. This could even result in an estimate with a larger MSE than the human-only baseline (for sensitivity to miscalibration, see App B.2). However, by using power tuning, as detailed in Section 3.2, we ensure that incorporating LLM annotation into the estimation process does not hurt the estimate (i.e., does not increase the MSE) regardless of the sampling method used for human annotations (whether uniform or adaptive).
@@
Finally, the accessibility of our method is an important consideration. Across disciplines, researchers can simply prompt the LLM for its confidence via API access and leverage Confidence-Driven Inference to combine LLM confidence with LLM and human annotations to produce a valid statistical estimate. This approach can be applied to a wide range of tasks, across fields.
Limitations
@@
Additionally, while we treat human annotations as the gold standard in our study, we acknowledge that human annotations are biased, and that reasonable annotators can disagree, making it necessary to account for annotator-specific parameters Hashemi et al. (2024). Future work could explore ways to account for variability and bias in human annotations.
Human annotations are often obtained through crowdsourcing, which may itself be influenced by LLMs, as crowd workers might use LLMs to increase productivity Veselovsky et al. (2023). Although we use datasets collected before the widespread availability of LLMs, detecting AI-generated text remains a challenge Verma et al. (2024).
@@
Acknowledgments
This work is supported by the Swiss National Science Foundation (Grant P500PT-211127), the Stanford Institute for Human-Centered Artificial Intelligence, Stanford Data Science, and Navy Grant N00014-24-1-2305.
@@ A.3 LLM and Human Annotation Details
Stage 1 GPT-4o annotations are in moderate agreement with human annotations in all three settings: $\kappa_{\mathrm{politeness}}=0.39$, $\kappa_{\mathrm{stance}}=0.57$, and $\kappa_{\mathrm{bias}}=0.43$. For context, human annotators had a median inter-annotator pairwise correlation of 0.68 for the politeness dataset, while average inter-annotator agreement ranged from 0.54 to 0.64 across annotation rounds for the stance datasetDanescu-Niculescu-Mizil et al. (2013); Luo et al. (2020). No agreement data is available for the political bias dataset.
@@ A.4 Computation of Evaluation Metrics
Appendix B Supplementary Results
B.1 LLM Data Collection Robustness
To examine the robustness of our evaluation to choices in the LLM data collection, we collected LLM annotations using varying approaches on the task of analyzing impact of hedging on perceived politeness. The default experiment (cf. Figure 2) leverages LLM annotations collected using GPT-4o model, via zero-shot prompting, where the annotation task is binary classification. In Figure 3, for $n_{human}=1100$, we report the gain in effective sample size and coverage using (1) an alternative smaller model (GPT-4o-mini), (2) alternative prompting mechanism (few shot prompting with ten examples), and (3) alternative annotation task (rating on a 7-point bipolar Likert scale, ranging from “very impolite (1)” to “very polite (7)”).
For each LLM data collection method, the confidence-driven approach consistently achieves a higher gain in effective sample size than the non-adaptive approach. Moreover, LLM-only coverage is poor across the different data collection methods (except for the experiment with a smaller model), while non-adaptive and adaptive approaches achieve 90% coverage or higher. We note that, although inter-annotator agreement varies substantially depending on these choices, between low ($\kappa_{\mathrm{politeness}}=0.21$) and moderate ($\kappa_{\mathrm{politeness}}=0.56$), confidence-driven approach is not harmed by the varying quality of the annotations, and always achieves a positive gain in the effective sample size. We also note that with ten few-shot examples, LLM-only coverage increases (82%), as predicted, since examples help to guide the annotation task. We also note that LLM annotations are lower in quality when collected using a Likert scale, likely due to eliciting more fine-grained classification, which makes the rating task more challenging.
Additionally, Figure 4 and Table 4 summarize the results using GPT-3.5. For $n_{\mathrm{human}}=500$, in each estimation task the confidence-driven approach again achieves a higher gain in effective sample size than the non-adaptive approach. Moreover, it always achieves a positive gain. In contrast, the non-adaptive approach achieves a negative gain in three out of the five estimation tasks (both politeness estimates and the stance estimate). The confidence-driven and non-adaptive approaches always achieve over 90% coverage. In contrast, LLM-only coverage is always poor using GPT-3.5 (while using GPT-4o it was poor on four out of the five estimation tasks).
In summary, our insights regarding the gains of the confidence-driven approach are robust to the choices made in the LLM data collection.
B.2 Sensitivity to Confidence Calibration
@@
Method Metric
Gain in eff. sample size Coverage
default experiment, cf. Figure 2
(model: GPT-4o, prompting: zero-shot, annotation task: binary classification)
Cohen inter-rater agreement $\kappa_{\mathrm{politeness}}=0.39$
LLM only 71%
human + LLM (non-adaptive) -9.85% 93%
confidence-driven 248% 99%
alternative model: GPT-4o-mini
Cohen inter-rater agreement $\kappa_{\mathrm{politeness}}=0.56$
LLM only 90%
human + LLM (non-adaptive) 24.15% 94%
confidence-driven 265.69% 100%
alternative prompting: few-shot (10 examples)
Cohen inter-rater agreement $\kappa_{\mathrm{politeness}}=0.45$
LLM only 82%
human + LLM (non-adaptive) 3.1% 91%
confidence-driven 251.25% 100%
alternative annotation task: Likert scale (7-point)
Cohen inter-rater agreement $\kappa_{\mathrm{politeness}}=0.21$
LLM only 26%
human + LLM (non-adaptive) -27.45% 92%
confidence-driven 239.92% 99%
Table 3: Sensitivity to the LLM data collection method. Gain in effective sample size and coverage for the LLM only, human + LLM (non-adaptive), and confidence-driven approaches, across varying data collection approaches. Results are presented for the task of analyzing impact of hedging on perceived politeness, $n_{\mathrm{human}}=1100$, estimated over 100 trials. The confidence-driven approach always achieves a large   positive gain. For each LLM data collection method, the confidence-driven approach achieves a higher gain in effective sample size than the non-adaptive approach (marked in bold); it also achieves near 90% coverage or higher in each setting. In contrast, LLM-only coverage is often poor. Gain in effective sample size is not estimated for the LLM-only approach as it does not leverage human annotations.
@@
Stance on global warming Gain in eff. sample size ( 82.68 $\pm$ 9.17)( -11.23 $\pm$ 17.15)%
Coverage 96% 94% 0%